A man is able to save N50.00 of his salary in a certain year and afterwards saves in any one year N20.00 more than he saved in the preceding year. how long does it take him to save N4370.00?
Answer by Boreal(15235)

sum=an=4370. This equals (n/2)(2a1+(n-1)d). a1=50 and d=20
so multiplying by 2, and 8740=n(100+20n-20)
100n+20n^2-20n=8740
20n^2+80n-8740=0
n^2+4n-437=0
n=-4+/-sqrt(1764)/2
=(1/2)(-4+42), positive root only
=19 years ANSWER
divide everything by 10
5/7/9/11/13/15/17/19/21/23/25/27/29/31/33/35/37/39/41=437